# An explanation of the number of points and symmetries of starbursts

### v1

@@ 1. Introduction
The starbursts are only perceived when the eye is affected by a noticeable amount of aberrations [21, 26]. These are mathematically described as the wave aberration function, defined as the difference between the actual wavefront and the ideal reference spherical wavefront that would eventually concentrate all rays in a single point. Considering that the angular size of a typical star is around the maximum resolution of the eye, namely, 1 arcmin, the image of a star at the retina can only be resolved if the amount of wave aberration generates caustics that exceed that threshold resolution value of 1 arcmin [16].
@@ 2.1. Symmetry properties of startbursts
The maximum number of saddle cusps of Gauss is equal to $(n-2)(2n-5)$, being $n$ the degree of the wave aberration function $W$, expressed as a linear combination of Zernike polynomials.
@@
If a surface described by a bivariate polynomial has only isolated critical points, then the number of isolated extrema is bounded above. Also, this bound depends upon the degree of the variables of the polynomial [19]. Let’s say $d$ is the degree of the bivariate polynomial $f$ containing only isolated critical points. Then [9] the number of isolated extrema, denoted by $e$, meets the condition: $e\leq 0.5(d^{2}-d+1)$. Moreover, the number of saddle points ($s$) is bounded to ([9, eq. (3.1)]): $s\leq 0.5(d^{2}-d)$.
Now, let us consider $W$ the wave aberration function, expressed as a linear combination of Zernike polynomials, being $n$ its degree. Thus, the degree of $G=\det({\rm Hess}(W))$ is: $\deg(G)\leq 2(n-2)$. Therefore, the number of saddle points of $G$ is, at most, $(n-2)(2n-5)$. ∎
@@
With this, we have proved that if $(\rho^{*},0)$ is a critical point, then $(\rho^{*},\frac{2k\pi}{n})$ for $k=0,...,n-1$ are also critical points.
@@
An analog argument, as in the previous case, leads us to the same conclusion: if $(\rho^{*},\pi)$ is critical point, then $(\rho^{*},\pi+\frac{2k\pi}{n})$ for $k=0,...,n-1$ are also critical points.
@@
Let us consider a wave aberration function defined as in (1), with $n\in\{3,4,5,6\}$. Then, the number $p$ or saddle cusps of Gauss inside the pupil is:
@@
    $p=0$, $p=n$ or $p=2n$ when $n=5,6$.
Moreover, if $p=n$, then the points are uniformly distributed in a circle centered at the origin. Also, if $p=2n$, they are uniformly distributed in two circles of different radius, $n$ points in each.
@@
Observe here that, if they are real, there are two positive values and two negative ones. To give conditions for these values to be real, $10<\rho<$ 1, and the corresponding critical points to be saddle cusps of Gauss, we need to define the following parameters:
@@
$\gamma$$\alpha$$\alpha_{3}$$\alpha_{1}^{+}$$\alpha_{2}$$\alpha_{1}^{-}$$\alpha_{3}$$\alpha_{1}^{+}$$\alpha_{2}$$\alpha_{1}^{-}$$4\sqrt{10}\beta$$-4\sqrt{10}\beta$$-2\sqrt{15}\beta$$\sqrt{15}\beta$$12\sqrt{15}\beta$$\bullet$
@@
    $\gamma>0$ $\alpha_{1}^{-}<\alpha<\alpha_{2}$ $-4\sqrt{10}\beta<\gamma<0$ $\alpha_{2}<\alpha<\alpha_{3}$ $\gamma<-4\sqrt{10}\beta$ $\alpha_{2}<\alpha<\alpha_{1}^{-}$
@@
$2\beta(\sqrt{15}\alpha-15\beta)+15\rho^{2}(6\beta^{2}-\gamma^{2}\pm 2\sqrt{2}\beta\gamma)=0,$ (7)
@@
[…]
@@
    $-3\sqrt{2}\beta<\gamma<0$ $\alpha_{1}^{+}<\alpha<\displaystyle\sqrt{15}\beta$ $\gamma>\sqrt{2}\beta$ $\sqrt{15}\beta<\alpha<\alpha_{1}^{+}$
@@
$\gamma$$\alpha$$\alpha_{1}^{+}$$\alpha_{1}^{-}$$\sqrt{15}\beta-\alpha_{2}^{+}$$\sqrt{15}\beta-\alpha_{2}^{-}$$-\gamma_{1}^{+}$$\gamma_{1}^{+}$$-\gamma_{1}^{-}$$\gamma_{1}^{-}$$2.26\beta$$-2.26\beta$$\sqrt{15}\beta$$\bullet$
@@
    $\gamma<-\gamma_{1}^{+}$ $\sqrt{15}\beta-\alpha_{2}^{+}<\alpha<\sqrt{15}\beta$ $-\gamma_{1}^{+}<\gamma<0$ $\alpha_{1}^{+}<\alpha<\sqrt{15}\beta$ $\gamma>\gamma_{1}^{-}$ $\sqrt{15}\beta-\alpha_{2}^{-}<\alpha<\alpha_{1}^{+}$
@@
$\gamma$$\alpha$$\alpha_{1}^{+}$$\alpha_{1}^{-}$$\sqrt{15}\beta+\alpha_{2}^{-}$$\sqrt{15}\beta-\alpha_{2}^{-}$$\sqrt{15}\beta+\alpha_{2}^{+}$$\sqrt{15}\beta-\alpha_{2}^{+}$$-\gamma_{1}^{+}$$-\gamma_{1}^{-}$$\gamma_{1}^{+}$$\gamma_{1}^{-}$$2.2\beta$$-2.2\beta$$\sqrt{15}\beta$$\bullet$
@@
    $\gamma>\gamma_{1}^{+}$ $\sqrt{15}\beta-\alpha_{2}^{-}<\alpha<\alpha_{1}^{+}$ $-\gamma_{1}^{+}<\gamma<0$ $\alpha_{1}^{+}<\alpha<\sqrt{15}\beta$ $\gamma\leqslant-\gamma_{1}^{+}$ $\sqrt{15}\beta+\alpha_{2}^{+}<\alpha<\sqrt{15}\beta$
@@ 2.2. Relation between Zernike coefficients: necessary conditions for the appearance of starbursts
Under low-light conditions (required to see starbursts), a typical value of pupil radius is $3.5$ $mm$; we use this value from now on. Some population statistical studies of Zernike aberration structure in normal and healthy eyes have revealed that higher-order aberrations tend to be randomly distributed about a mean value of zero [24], except for spherical aberration. Also, there is no apparent correlation between the degree of ametropia and spherical aberration [11]. The mean value of spherical aberration ($Z_{4}^{0}$ accompanying coefficient in the Zernike expansion of the wave aberration function) is positive and around $0.18\,\mu m$ over a $3\,mm$ pupil radius ($Rp$[11]. Taking into account these reference values, in (1) we set $\beta=0.2\,\mu m$ using a $3.5\,mm$ pupil radius. The coefficients accompanying the Zernike terms are given in $\mu m$. Also, as a gross approximation to spherical equivalent from the defocus Zernike coefficient ($\alpha$ in (1)) is given by formula [24]: $M=\frac{4\sqrt{3}\,\alpha}{Rp^{2}}$, being $R_{p}$ the pupil radius.
@@ 3.2. Arquetypical examples of starbursts
As previously mentioned, we concentrate on star-like starbursts, i.e., with a $p$-fold symmetry. A $p$-fold symmetric starburst can be classified depending on the number of points. A first classification is based on the parity of the number of points, so that we can say an odd or even starburst. A second classification concerns the number of points and their length. There are two possibilities: $p$ equally-spaced points from the starburst center, or $2p$, where $p$ points are the same lenght, and the other $p$ points have another lenght; one set is due to cusp caustics and the other to two very close fold caustics (see hypothesis 1). We name the first equally-distanced points starbursts and the second non-equally-distanced points starbursts. Our methodology allowed us to find four archetypical starbursts: $3$, $4$, $5$, and $6$ equally-distanced points starbursts, and an $8$ non-equally-distanced points starburst.
@@ 3.2.2. Even equally-distanced, and non-equal-distanced points starbursts
Now, the number of cusps of Gauss is $7$, $6$ of which are fertile, configuring $6$ lines of reflectional symmetry that determine a $6$-fold caustic pattern symmetry. As in the case of the 5-point staburst, each fertile cusp of Gauss generates a cusp caustic, and the action of two opposite fertile points, two converging fold caustics; the combination of both generates a long-bright starbust point. Thus inducing an equally distant 6-point staburst.

### v2

@@ 1. Introduction
The starbursts are only perceived when the eye is affected by a noticeable amount of aberrations [21, 26]. These are mathematically described as the wave aberration function, defined as the difference between the actual wavefront and the ideal reference spherical wavefront that would eventually concentrate all rays in a single point. Considering that the angular size of a typical star is much smaller than the maximum resolution of the eye, namely, $1$ arcmin, the image of a star at the retina can only be resolved if the amount of wave aberration generates caustics that exceed that threshold resolution value of $1$ arcmin [16].
@@ 2.1. Symmetry properties of startbursts
The maximum number of saddle cusps of Gauss is equal to (n2)(2n5)(n-2)(2n-5), being $n$ the degree of the wave aberration function $W$, expressed as a linear combination of Zernike polynomials.
@@
If a surface described by a bivariate polynomial has only isolated critical points, then the number of isolated extrema is bounded above. Also, this bound depends upon the degree of the variables of the polynomial [19]. Let’s say $d$ is the degree of the bivariate polynomial $f$ containing only isolated critical points. Then, the number of saddle points is bounded above by $0.5(d^{2}-d)$ [9, eq. (3.1)].
Now, let us consider $W$ the wave aberration function, expressed as a linear combination of Zernike polynomials, being $n$ its degree. Thus, the degree of $G=\det({\rm Hess}(W))$ is: $\deg(G)\leq 2(n-2)$. Therefore, the number of saddle points of $G$ is, at most, (n2)(2n5)(n-2)(2n-5). ∎
@@
With this, we have proved that if $(\rho^{*},0)$ is a critical point, then $(\rho^{*},\frac{2k\pi}{n})$ for k=0,,n1k=0,...,n-1 are also critical points.
@@
An analog argument, as in the previous case, leads us to the same conclusion: if $(\rho^{*},\pi)$ is critical point, then $(\rho^{*},\pi+\frac{2k\pi}{n})$ for k=0,,n1k=0,...,n-1 are also critical points.
@@
Let us consider a wave aberration function defined as in (1), with $n\in\{3,4,5,6\}$. Then, the number $s$ of saddle cusps of Gauss inside the pupil is:
@@
    $s=0$, $s=n$ or $s=2n$ when $n=5,6$.
Moreover, if $s=n$, then the points are uniformly distributed in a circle centered at the origin. Also, if $s=2n$, they are uniformly distributed in two circles of different radius, $n$ points in each.
@@
Observe here that, if they are real, there are two positive values and two negative ones. To give conditions for these values to be real, $0<\rho<1$, and the corresponding critical points to be saddle cusps of Gauss, we need to define the following parameters:
@@
$\gamma$$\alpha$$\alpha_{3}$$\alpha_{1}^{+}$$\alpha_{2}$$\alpha_{1}^{-}$$\alpha_{3}$$\alpha_{1}^{+}$$\alpha_{2}$$\alpha_{1}^{-}$$4\sqrt{10}\beta$$4\sqrt{10}\beta$$2\sqrt{15}\beta$$\sqrt{15}\beta$$12\sqrt{15}\beta\bullet$
@@
    $\gamma>0$ $\alpha_{1}^{-}<\alpha<\alpha_{2}$ $4\sqrt{10}\beta<\gamma<0$ $\alpha_{2}<\alpha<\alpha_{3}$ $\gamma<-4\sqrt{10}\beta$ $\alpha_{2}<\alpha<\alpha_{1}^{-}$
@@
[…]
@@
$\gamma$$\alpha$$\alpha_{1}^{+}$$\alpha_{1}^{-}$$\sqrt{2}\beta$$3\sqrt{2}\beta$$\sqrt{2}\beta$$3\sqrt{2}\beta$     $3.8\beta$$3.8\beta$  $\sqrt{15}\beta\bullet\diamond$
@@
    $3\sqrt{2}\beta<\gamma<0$ $\alpha_{1}^{+}<\alpha<\displaystyle\sqrt{15}\beta$ $\gamma>\sqrt{2}\beta$ $\sqrt{15}\beta<\alpha<\alpha_{1}^{+}$
@@
$\gamma$$\alpha$$\alpha_{1}^{+}$$\alpha_{1}^{-}$$\sqrt{15}\beta-\alpha_{2}^{+}$$\sqrt{15}\beta-\alpha_{2}^{-}$$\gamma_{1}^{+}$$\gamma_{1}^{+}$$\gamma_{1}^{-}$$\gamma_{1}^{-}$$26\beta2.2$$26\beta$$\sqrt{15}\beta\bullet$
@@
    $\gamma<-\gamma_{1}^{+}$ $\sqrt{15}\beta-\alpha_{2}^{+}<\alpha<\sqrt{15}\beta$ $\gamma_{1}^{+}<\gamma<0$ $\alpha_{1}^{+}<\alpha<\sqrt{15}\beta$ $\gamma>\gamma_{1}^{-}$ $\sqrt{15}\beta-\alpha_{2}^{-}<\alpha<\alpha_{1}^{+}$
@@
$\gamma$$\alpha$$\alpha_{1}^{+}$$\alpha_{1}^{-}$$\sqrt{15}\beta+\alpha_{2}^{-}$$\sqrt{15}\beta-\alpha_{2}^{-}$$\sqrt{15}\beta+\alpha_{2}^{+}$$\sqrt{15}\beta-\alpha_{2}^{+}$$\gamma_{1}^{+}$$\gamma_{1}^{-}$$\gamma_{1}^{+}$$\gamma_{1}^{-}$$2\beta2.2$$2\beta$$\sqrt{15}\beta\bullet$
@@
    $\gamma>\gamma_{1}^{+}$ $\sqrt{15}\beta-\alpha_{2}^{-}<\alpha<\alpha_{1}^{+}$ $\gamma_{1}^{+}<\gamma<0$ $\alpha_{1}^{+}<\alpha<\sqrt{15}\beta$ $\gamma\leqslant-\gamma_{1}^{+}$ $\sqrt{15}\beta+\alpha_{2}^{+}<\alpha<\sqrt{15}\beta$
@@ 2.2. Relation between Zernike coefficients: necessary conditions for the appearance of starbursts
Under low-light conditions (required to see starbursts), a typical value of pupil radius is $3.5$ mmmm; we use this value from now on. Some population statistical studies of Zernike aberration structure in normal and healthy eyes have revealed that higher-order aberrations tend to be randomly distributed about a mean value of zero [24], except for spherical aberration. Also, there is no apparent correlation between the degree of ametropia and spherical aberration [11]. The mean value of spherical aberration ($Z_{4}^{0}$ accompanying coefficient in the Zernike expansion of the wave aberration function) is positive and around $0.18\,\mu m$ over a 3mm3\,$m$ pupil radius (RpRp[11]. Taking into account these reference values, in (1) we set $\beta=0.2\,\mu m$ using a 3.5mm3.5\,$m$ pupil radius. The coefficients accompanying the Zernike terms are given in $\mu m$. Also, as a gross approximation to spherical equivalent from the defocus Zernike coefficient ($\alpha$ in (1)) is given by formula [24]: $M=\frac{4\sqrt{3}\,\alpha}{Rp^{2}}$, being $R_{p}$ the pupil radius.
@@ 3.2. Arquetypical examples of starbursts
As previously mentioned, we concentrate on star-like starbursts, i.e., with a $p$-fold symmetry. A $p$-fold symmetric starburst can be classified depending on the number of points. A first classification is based on the parity of the number of points, so that we can say an odd or even starburst. A second classification concerns the number of points and their length. There are two possibilities: $p$ equally-spaced points from the starburst center, or $2p$, where $p$ points are the same lenght, and the other $p$ points have another lenght; one set is due to cusp caustics and the other to two very close fold caustics (see hypothesis 1). We name the first equally-distanced points starbursts and the second non-equally-distanced points starbursts. Our methodology allowed us to find four archetypical starbursts: $3$, 44, 55, and $6$ equally-distanced points starbursts, and an $8$ non-equally-distanced points starburst.
@@ 3.2.2. Even equally-distanced, and non-equal-distanced points starbursts
Now, the number of cusps of Gauss is $7$, 66 of which are fertile, configuring $6$ lines of reflectional symmetry that determine a $6$-fold caustic pattern symmetry. As in the case of the 5-point staburst, each fertile cusp of Gauss generates a cusp caustic, and the action of two opposite fertile points, two converging fold caustics; the combination of both generates a long-bright starbust point. Thus inducing an equally distant 6-point staburst.